# Voltage-sensitive distribution factors for contingency analysis and topology optimization

### v1

@@ Abstract
Topology optimization is a promising approach for mitigating congestion and managing changing grid conditions, but it is computationally challenging and requires approximations. Conventional distribution factors like PTDFs and LODFs, based on DC power flow, fail to capture voltage variations, reactive power, and losses — limiting their use in detailed optimization tasks such as busbar splitting. This paper introduces generalized distribution factors derived from a voltage-sensitive linearization of the full AC power flow equations. The proposed formulation accurately reflects reactive power flows, Ohmic losses, and voltage deviations while remaining computationally efficient. We derive and evaluate generalized PTDFs, LODFs, and topology modification factors using matrix identities. We discuss potential applications including voltage-aware N1N-1 security analysis, and topology optimization with a focus on busbar splitting. Numerical experiments demonstrate close agreement with full AC solutions, significantly outperforming the traditional DC approximation.
Index terms– Bus Splitting, Line Outage Distribution Factor, Linearized AC model, Power Transfer Distribution Factor, Topology Optimization
@@ I Introduction
Distribution factors are standard tools in the design and operations of transmission grids. The standard power transfer distribution factors (PTDFs) and line outage distribution factors (LODFs) describe how line flows change when power injections are adapted or transmission lines fail [9]. Recently, distribution factors for other topology changes have been introduced for grid planning and topology optimization [10, 11]. All these distribution factors are derived using the DC approximation, a linearized form of the power flow equations [9], which is valid when voltages are close to their set points and losses and grid loads are moderate [12, 13].
In recent years, several generalizations of the DC approximation have been proposed that avoid some of the assumptions and simplifications, see Sec. II for a literature review. The derivation generally starts from the full nonlinear load-flow equations, which include reactive power and Ohmic losses. In this article, we derive distribution factors for generalized linear power flow approximations including reactive power, losses and voltage variations. We provide expressions generalizing PTDFs, LODFs and bus merge and split distribution factors.
@@ II Literature Review
Several generalizations that try to avoid these assumptions have been developed: (i) Models that incorporate Ohmic losses but keep the assumption of fixed voltage magnitudes have been proposed for instance in Refs. [13, 16]. (ii) Methods to reduce or bound the linearization error [17, 16, 15]. (iii) Several approaches that linearize the full AC load-flow equations, keeping Ohmic losses and voltage variations, have been proposed. This is the focus of the current manuscript. To arrive at a linear set of equations, most authors introduce the approximations,
@@
III Linearized load-flow equations
III-A Linearizing the AC load-flow equations
@@
We can follow two strategies for the definition of the reference state: (i) For general computations, we can use a flat initial guess where $\hat{\theta}_{ft}=0$ for ordinary transmission lines and $\hat{\theta}_{ft}=\alpha_{ft}$ for PSTs. This strategy is commonly referred to as a cold-start model. (2) To assess line outages and topology modifications, we can use the pre-modification state for $\hat{\theta}_{ft}$. This strategy is commonly referred to as a hot-start model.
@@ III-B Shunt elements
$\begin{pmatrix}p_{f}\\ q_{f}\end{pmatrix}=\begin{pmatrix}{\hat{p}}_{f}^{\text{shunt}}\\ \hat{q}_{f}^{\text{shunt}}\end{pmatrix}+\begin{pmatrix}N^{p\theta}_{f}&N^{p\theta}_{f}\\ N^{pu}_{f}&N^{pu}_{f}\\ \end{pmatrix}\begin{pmatrix}\theta_{f}\\ u_{f}\end{pmatrix}$ (3)
@@ III-C The network equations
We now collect all prior results to formulate the linearized load-flow equations for an extended grid. We number the PQ nodes as $i=1,\ldots,n$, PV nodes as $i=n+1,\ldots,n+m$ and the slack node as $n+m+1$. The total number of nodes is denoted as $\mathtt{N}=n+m+1$.
@@
We note that the matrix $\boldsymbol{M}$} corresponds to the Jacobian of the AC load-flow equations at the reference state. We thus obtain the linearized load flow equations in a compact vectorial form
@@ IV-A Branch modification distribution factors
We start from a reference grid for which the linearized power flow equations are given by Eq. (6). Notably, if we have solved the full nonlinear load-flow equations for this reference grid, we may use the resulting voltage magnitudes as the reference values $\hat{v}_{i}$ and $\hat{\theta}_{i}$.
@@
To compute $\boldsymbol{\psi}$} and $\boldsymbol{\sigma}$}, we subtract the linear power flow equations (6) before and after the branch modification and obtain
$\begin{pmatrix}\Delta{\boldsymbol{\hat{p}}}\\ \Delta{\boldsymbol{\hat{q}}}\end{pmatrix}+\Delta{\boldsymbol{M}}\begin{pmatrix}{\boldsymbol{\theta}}\\ {\boldsymbol{u}}\end{pmatrix}=-\left({\boldsymbol{M}}+\Delta{\boldsymbol{M}}\right)\begin{pmatrix}{\boldsymbol{\psi}}\\ {\boldsymbol{\sigma}}\end{pmatrix}$ (8)
@@
$={\boldsymbol{M}}^{-1}-{\boldsymbol{M}}^{-1}{\boldsymbol{S}}\left(\mathds{1}+{\boldsymbol{R}}{\boldsymbol{M}}^{-1}{\boldsymbol{S}}\right)^{-1}{\boldsymbol{R}}{\boldsymbol{M}}^{-1},$ (10)
where $\mathds{1}$ denotes the identity matrix. We emphasize that the matrix $\left(\mathds{1}+{\boldsymbol{R}}{\boldsymbol{M}}^{-1}{\boldsymbol{S}}\right)$ has size $3\times 3$, such that we can easily compute its inverse. This approach is particularly important if we want to evaluate various branch modifications, as for instance all transmission line outages for a comprehensive security assessment. In this case we have to invert $\boldsymbol{M}$} only once for the reference grid topology. For any branch modification we can then compute $\left({\boldsymbol{M}}+\Delta{\boldsymbol{M}}\right)^{-1}$ efficiently using the Woodbury matrix identity.
@@ IV-B Multiple branch modifications
and is of rank $3k$. The change in the state variables is still given by Eq. (9), and the inverse can be computed using the Woodbury matrix identity in Eq. (10), where the matrix $\left(\mathds{1}+{\boldsymbol{R}}{\boldsymbol{M}}^{-1}{\boldsymbol{S}}\right)$ now has size $3k\times 3k$.
@@ IV-C Line modification and outage distribution factors
$LMDF_{(kl),(ij)}$ $={\boldsymbol{A}}\left(\mathds{1}+{\boldsymbol{D}}{\boldsymbol{A}}\right)^{-1}{\boldsymbol{D}}$ (13)
@@ IV-D Numerical evaluation
We test the accuracy of the generalized DC+ approximation by comparing to full AC simulations. We use plain AC load-flow computations in PyPowSyBl [33] without any adaptable grid elements. That is, we disable any type of controller, generator limits, distributed slack and DC grid elements and ensure that the $\pi$-line model is used for all transformers.
In each test we first carry out an AC load-flow for the N0N-0 case and use the results for the reference values $\hat{v}_{i}$ and $\hat{\theta}_{ft}$ and the computation of the matrix $\boldsymbol{M}$}. Then we run an N1N-1 analysis, simulating the failure of each branch in the grid. For each failure, we compute the state variables $v_{m}$, $\theta_{m}$ and the power injections $P_{m}$ and $Q_{m}$ for each bus $i$ via the generalized distribution factors given by Eq. (9). We then evaluate the approximation errors, i.e. the deviation from to the full AC results. In addition, we compute the approximation error for the common DC approximation for comparison.
@@
A detailed analysis for the IEEE 14-bus test grid is provided in Fig. 1. We find that the approximation error in the state variables is very small for almost all outage scenarios; with moderate errors for the outage of branch 1. Most importantly, the generalized DC+ approximation given by Eq. (9) captures the changes in voltage magnitudes well. Errors are most pronounced in the nodal reactive power injections.
@@ V Switches and busbar couplers
Dynamic topology reconfiguration is a promising option for reducing grid congestion [6]. This includes, in particular, opening or closing switches or busbar couplers. In the following, we will assume that the inverse of the matrix, ${\boldsymbol{M}}^{-1}$, has been computed once for a reference topology. We show how to efficiently update this matrix after a switch or coupler is modified, to support the numerical optimization of topology reconfiguration.
@@ V-A Closing of switches and busbar couplers
We begin by treating the closing of a switch using methods from Sec. IV-A. We first assume that $b_{e}^{s}$ is finite and take the limit $b_{e}^{s}\rightarrow\infty$ only in the final result. In the following, the symbol ${\boldsymbol{M}}_{o}$ denotes the grid with the switch open and ${\boldsymbol{M}}_{c}$ the grid with the switch closed. We thus have
${\boldsymbol{M}}_{c}$ $={\boldsymbol{M}}_{o}+\Delta{\boldsymbol{M}},$
@@
$=$ ${\boldsymbol{M}}_{o}^{-1}-{\boldsymbol{M}}_{o}^{-1}{\boldsymbol{S}}\left[\frac{1}{b_{e}^{s}\hat{v}_{f}\hat{v}_{t}}\mathds{1}+{\boldsymbol{R}}{\boldsymbol{M}}_{o}^{-1}{\boldsymbol{S}}\right]^{-1}{\boldsymbol{R}}{\boldsymbol{M}}_{o}^{-1}$
@@
${\boldsymbol{M}}_{c}^{-1}=$ ${\boldsymbol{M}}_{o}^{-1}-{\boldsymbol{M}}_{o}^{-1}{\boldsymbol{S}}\left[{\boldsymbol{R}}{\boldsymbol{M}}_{o}^{-1}{\boldsymbol{S}}\right]^{-1}{\boldsymbol{R}}{\boldsymbol{M}}_{o}^{-1}$ (17)
We can now compute the state variables for the grid with the closed switch by solving Eq. (6) and carry out an N1N-1 contingency analysis using Eq. (9) – using ${\boldsymbol{M}}_{c}^{-1}$ instead of ${\boldsymbol{M}}_{o}^{-1}$.
@@ V-B Bus splits
In practice, bus splits are often more important than bus mergers because busbar couplers are typically closed in the reference grid layout. In this section we introduce a formulation of bus splits for a single busbar coupler. Initially, the coupler is closed such that the two busbars effectively constitute a single node of the network. We want to describe the impact of the opening of the coupler making the two busbars two separate nodes with no direct connection. For simplicity we assume that the two terminal busbars are treated as PQ nodes without shunt elements.
@@
Let the matrix ${\boldsymbol{M}}_{m}$ describe the reference grid where all switches or busbar couplers are closed and merged into a single nodes. Furthermore, let ${\boldsymbol{M}}_{o}$ describe the grid where the busbar coupler has been opened. The two uncoupled busbars then constitute separate nodes so that the dimension of ${\boldsymbol{M}}_{o}$ is larger than the dimension of ${\boldsymbol{M}}_{m}$. To relate the case of interest to the reference case, we introduce a third network configuration ${\boldsymbol{M}}_{c}$ where all switches and busbar couplers are closed, but where the nodes have not been merged. In this configuration, the incident branches must be assigned to the node (busbar), to which they are connected when the coupler opens. The two matrices ${\boldsymbol{M}}_{c}$ and ${\boldsymbol{M}}_{o}$ have the same dimension. These configurations are sketched in Fig. 3.
@@
$\Delta{\boldsymbol{M}}$ $=+b_{e}^{s}\hat{v}_{f}^{2}({\boldsymbol{\mu}}_{e}{\boldsymbol{\mu}}_{e}^{\top}+{\boldsymbol{\nu}}_{e}{\boldsymbol{\nu}}_{e}^{\top})$ (18)
@@
$\left[{\boldsymbol{M}}_{c}+\Delta{\boldsymbol{M}}\right]^{-1}={\boldsymbol{M}}_{c}^{-1}-{\boldsymbol{M}}_{c}^{-1}\begin{pmatrix}{\boldsymbol{\mu}}_{e}&{\boldsymbol{\nu}}_{e}\end{pmatrix}$
$\quad\times\left[\frac{1}{b_{e}^{s}\hat{v}_{f}^{2}}\mathds{1}+\begin{pmatrix}{\boldsymbol{\mu}}_{e}^{\top}\\ {\boldsymbol{\nu}}_{e}^{\top}\end{pmatrix}{\boldsymbol{M}}_{c}^{-1}\begin{pmatrix}{\boldsymbol{\mu}}_{e}&{\boldsymbol{\nu}}_{e}\end{pmatrix}\right]^{-1}\begin{pmatrix}{\boldsymbol{\mu}}_{e}^{\top}\\ {\boldsymbol{\nu}}_{e}^{\top}\end{pmatrix}{\boldsymbol{M}}_{c}^{-1}.$
@@
$(\mathds{1}-{\boldsymbol{M}}_{c}^{-1}{\boldsymbol{M}}_{o}){\boldsymbol{\mu}}_{e}=-{\boldsymbol{M}}_{c}^{-1}\,\Delta{\boldsymbol{M}}{\boldsymbol{\mu}}_{e}=-b_{e}^{s}\hat{v}_{f}^{2}{\boldsymbol{M}}_{c}^{-1}{\boldsymbol{\mu}}_{e}$
and an equivalent relation for ${\boldsymbol{\nu}}_{e}$. Using these relation as well as their transpose, we obtain
$\left[\frac{1}{b_{e}^{s}\,\hat{v}_{f}^{2}}\mathds{1}+\begin{pmatrix}{\boldsymbol{\mu}}_{e}^{\top}\\ {\boldsymbol{\nu}}_{e}^{\top}\end{pmatrix}{\boldsymbol{M}}_{c}^{-1}\begin{pmatrix}{\boldsymbol{\mu}}_{e}&{\boldsymbol{\nu}}_{e}\end{pmatrix}\right$] (19)
@@
${\boldsymbol{M}}_{o}^{-1}$ $={\boldsymbol{M}}_{c}^{-1}+\left(\mathds{1}-{\boldsymbol{M}}_{c}^{-1}{\boldsymbol{M}}_{o}\right)\begin{pmatrix}{\boldsymbol{\mu}}_{e}&{\boldsymbol{\nu}}_{e}\end{pmatrix}$
@@
$\quad\times\begin{pmatrix}{\boldsymbol{\mu}}_{e}^{\top}\\ {\boldsymbol{\nu}}_{e}^{\top}\end{pmatrix}\left(\mathds{1}-{\boldsymbol{M}}_{o}{\boldsymbol{M}}_{c}^{-1}\right).$ (20)

### v2

@@ Abstract
Topology optimization is a promising approach for mitigating congestion and managing changing grid conditions, but it is computationally challenging and requires approximations. Conventional distribution factors like PTDFs and LODFs, based on DC power flow, fail to capture voltage variations, reactive power, and losses, thereby limiting their use in detailed optimization tasks such as busbar splitting. This paper introduces generalized distribution factors derived from a voltage-sensitive linearization of the full AC power flow equations. The proposed formulation accurately reflects reactive power flows, Ohmic losses, and voltage deviations while remaining computationally efficient. We derive and evaluate generalized PTDFs, LODFs, and topology modification factors using matrix identities. We discuss potential applications including voltage-aware N1N-1 security analysis and topology optimization with a focus on busbar splitting. Numerical experiments demonstrate close agreement with full AC solutions, significantly outperforming the traditional DC approximation.
@@ I Introduction
Distribution factors are standard tools in the design and operation of transmission grids. The standard power transfer distribution factors (PTDFs) and line outage distribution factors (LODFs) describe how line flows change when power injections are adapted or transmission lines fail [9]. Recently, distribution factors for other topology changes have been introduced for grid planning and topology optimization [10, 11]. All these distribution factors are derived using the DC approximation, a linearized form of the power flow equations [9], which is valid when voltages are close to their set points and losses and grid loads are moderate [12, 13].
In recent years, several generalizations of the DC approximation have been proposed that avoid some of the assumptions and simplifications, see Sec. II for a literature review. The derivation generally starts from the full nonlinear power flow equations, which include reactive power and Ohmic losses. In this article, we derive distribution factors for generalized linear power flow approximations including reactive power, losses and voltage variations. We provide expressions generalizing PTDFs, LODFs and bus merge and split distribution factors.
An additional advantage of the proposed formulation is that it retains the algebraic structure of classical distribution factors, enabling efficient low-rank updates and highly parallel implementations. In particular, the computations map naturally to GPU architectures, allowing extremely high evaluation rates for contingency screening and topology analysis.
The article is organized as follows. Sec. II reviews related work on generalized linear power flow approximations. In Sec. III we derive the linearized power flow equations and introduce a convenient vectorial notation. Sec. IV develops generalized distribution factors for branch modifications and line outages. In Sec. V we extend the approach to switches and busbar couplers. The numerical performance is evaluated in Sec. VI before concluding in Sec. VII.
@@ II Literature Review
Several generalizations that try to avoid these assumptions have been developed: (i) Models that incorporate Ohmic losses but keep the assumption of fixed voltage magnitudes have been proposed for instance in Refs. [13, 16]. (ii) Methods to reduce or bound the linearization error [17, 16, 15]. (iii) Several approaches that linearize the full AC power flow equations, keeping Ohmic losses and voltage variations, have been proposed. This is the focus of the current manuscript. To arrive at a linear set of equations, most authors introduce the approximations,
@@
III Linearized power flow equations
III-A Linearizing the AC power flow equations
@@
We can follow two strategies for the definition of the reference state: (i) For general computations, we can use a flat initial guess where $\hat{\theta}_{ft}=0$ for ordinary transmission lines and $\hat{\theta}_{ft}=\alpha_{ft}$ for PSTs. This strategy is commonly referred to as a cold-start model. (ii) To assess line outages and topology modifications, we can use the pre-modification state for $\hat{\theta}_{ft}$. This strategy is commonly referred to as a hot-start model.
@@ III-B Shunt elements
$\begin{pmatrix}p_{f}\\ q_{f}\end{pmatrix}=\begin{pmatrix}{\hat{p}}_{f}^{\text{shunt}}\\ \hat{q}_{f}^{\text{shunt}}\end{pmatrix}+\begin{pmatrix}N^{p\theta}_{f}&N^{pu}_{f}\\ N^{q\theta}_{f}&N^{qu}_{f}\\ \end{pmatrix}\begin{pmatrix}\theta_{f}\\ u_{f}\end{pmatrix}$ (3)
@@ III-C The network equations
We now collect all prior results to formulate the linearized power flow equations for an extended grid. We number the PQ nodes as $i=1,\ldots,n$, PV nodes as $i=n+1,\ldots,n+m$ and the slack node as $n+m+1$. The total number of nodes is denoted as $\mathtt{N}=n+m+1$.
@@
We note that the matrix $\boldsymbol{M}$} corresponds to the Jacobian of the AC power flow equations at the reference state. We thus obtain the linearized power flow equations in a compact vectorial form
@@ IV-A Branch modification distribution factors
We start from a reference grid for which the linearized power flow equations are given by Eq. (6). Notably, if we have solved the full nonlinear power flow equations for this reference grid, we may use the resulting voltage magnitudes as the reference values $\hat{v}_{i}$ and $\hat{\theta}_{i}$.
@@
To compute $\boldsymbol{\psi}$} and $\boldsymbol{\sigma}$}, we subtract the linear power flow equations (6) before and after the modification and obtain
$\begin{pmatrix}\Delta{\boldsymbol{\hat{p}}}\\ \Delta{\boldsymbol{\hat{q}}}\end{pmatrix}+\Delta{\boldsymbol{M}}\begin{pmatrix}{\boldsymbol{\theta}}\\ {\boldsymbol{u}}\end{pmatrix}=-\left({\boldsymbol{M}}+\Delta{\boldsymbol{M}}\right)\begin{pmatrix}{\boldsymbol{\psi}}\\ {\boldsymbol{\sigma}}\end{pmatrix}$, (8)
@@
$={\boldsymbol{M}}^{-1}-{\boldsymbol{M}}^{-1}{\boldsymbol{S}}\left({\boldsymbol{I}}+{\boldsymbol{R}}{\boldsymbol{M}}^{-1}{\boldsymbol{S}}\right)^{-1}{\boldsymbol{R}}{\boldsymbol{M}}^{-1},$ (10)
where $\boldsymbol{I}$} denotes the identity matrix. We emphasize that the matrix $\left({\boldsymbol{I}}+{\boldsymbol{R}}{\boldsymbol{M}}^{-1}{\boldsymbol{S}}\right)$ has size $3\times 3$, such that we can easily compute its inverse. This approach is particularly important if we want to evaluate various branch modifications, as for instance all transmission line outages for a comprehensive security assessment. In this case we have to invert $\boldsymbol{M}$} only once for the reference grid topology. For any branch modification we can then compute $\left({\boldsymbol{M}}+\Delta{\boldsymbol{M}}\right)^{-1}$ efficiently using the Woodbury matrix identity.
@@ IV-B Multiple branch modifications
and is of rank $3k$. The change in the state variables is still given by Eq. (9), and the inverse can be computed using the Woodbury matrix identity in Eq. (10), where the matrix $\left({\boldsymbol{I}}+{\boldsymbol{R}}{\boldsymbol{M}}^{-1}{\boldsymbol{S}}\right)$ now has size $3k\times 3k$.
@@ IV-C Line modification and outage distribution factors
$LMDF_{(kl),(ij)}$ $={\boldsymbol{A}}\left({\boldsymbol{I}}+{\boldsymbol{D}}{\boldsymbol{A}}\right)^{-1}{\boldsymbol{D}}$ (13)
@@ IV-D Numerical evaluation
We test the accuracy of the generalized DC+ approximation by comparing it to full AC simulations. We use plain AC power flow computations in PyPowSyBl [33] without any adaptable grid elements. That is, we disable any type of controller, generator limits, distributed slack and DC grid elements and ensure that the $\pi$-line model is used for all transformers.
In each test we first carry out an AC power flow for the N0N-0 case and use the results for the reference values $\hat{v}_{i}$ and $\hat{\theta}_{ft}$ and the computation of the matrix $\boldsymbol{M}$}. Then we run an N1N-1 analysis, simulating the failure of each branch in the grid. For each failure, we compute the state variables $v_{m}$, $\theta_{m}$ and the power injections $P_{m}$ and $Q_{m}$ for each bus $i$ via the generalized distribution factors given by Eq. (9). We then evaluate the approximation errors, i.e. the deviation from the full AC results. In addition, we compute the approximation error for the common DC approximation for comparison.
@@
A detailed analysis for the IEEE 14-bus test grid is provided in Fig. 1. We find that the approximation error in the state variables is very small for almost all outage scenarios; with moderate errors for the outage of branch 1. Most importantly, the generalized DC+ approximation given by Eq. (9) captures the changes in voltage magnitudes well. Errors are most pronounced for nodal reactive power injections.
@@ V Switches and busbar couplers
Dynamic topology reconfiguration is a promising option for reducing grid congestion [6]. This includes, in particular, opening or closing switches or busbar couplers. In the following, we will assume that the inverse of the matrix, ${\boldsymbol{M}}^{-1}$, has been computed once for a reference topology. We show how to efficiently update this matrix after a switch or busbar coupler is modified.
@@ V-A Closing of switches and busbar couplers
We base our analysis on the results of Sec. IV-A. We first assume that $b_{e}^{s}$ is finite and take the limit $b_{e}^{s}\rightarrow\infty$ only in the final result. In the following, the symbol ${\boldsymbol{M}}_{o}$ denotes the grid with the switch open and ${\boldsymbol{M}}_{c}$ the grid with the switch closed. We thus have ${\boldsymbol{M}}_{c}={\boldsymbol{M}}_{o}+\Delta{\boldsymbol{M}}$ with
@@
$=$ ${\boldsymbol{M}}_{o}^{-1}-{\boldsymbol{M}}_{o}^{-1}{\boldsymbol{S}}\left[\frac{1}{b_{e}^{s}\hat{v}_{f}\hat{v}_{t}}{\boldsymbol{I}}+{\boldsymbol{R}}{\boldsymbol{M}}_{o}^{-1}{\boldsymbol{S}}\right]^{-1}{\boldsymbol{R}}{\boldsymbol{M}}_{o}^{-1}$
@@
${\boldsymbol{M}}_{c}^{-1}=$ ${\boldsymbol{M}}_{o}^{-1}-{\boldsymbol{M}}_{o}^{-1}{\boldsymbol{S}}\left[{\boldsymbol{R}}{\boldsymbol{M}}_{o}^{-1}{\boldsymbol{S}}\right]^{-1}{\boldsymbol{R}}{\boldsymbol{M}}_{o}^{-1}\,.$ (17)
We can now compute the state variables for the grid with the closed switch by solving Eq. (6) and carry out an N1N-1 contingency analysis via Eq. (9) using ${\boldsymbol{M}}_{c}^{-1}$ instead of ${\boldsymbol{M}}_{o}^{-1}$.
@@ V-B Bus splits
Bus splits are typically more important than bus mergers because busbar couplers are typically closed in the reference grid layout. In the following, we introduce a formulation of bus splits for a single busbar coupler. Initially, the coupler is closed such that the two busbars effectively constitute a single node of the network. We want to describe the impact of the opening of the coupler making the two busbars two separate nodes with no direct connection. For simplicity we treat the two terminal busbars as PQ nodes without shunt elements.
@@
Let the matrix ${\boldsymbol{M}}_{m}$ describe the reference grid where all switches or busbar couplers are closed and merged into single nodes. Furthermore, let ${\boldsymbol{M}}_{o}$ describe the grid where the busbar coupler has been opened. The two uncoupled busbars then constitute separate nodes so that the dimension of ${\boldsymbol{M}}_{o}$ is larger than the dimension of ${\boldsymbol{M}}_{m}$. To relate the case of interest to the reference case, we introduce a third network configuration ${\boldsymbol{M}}_{c}$ where all switches and busbar couplers are closed, but where the nodes have not been merged. In this configuration, the incident branches must be assigned to the node (busbar), to which they are connected when the coupler opens. The two matrices ${\boldsymbol{M}}_{c}$ and ${\boldsymbol{M}}_{o}$ have the same dimension. These configurations are sketched in Fig. 3.
@@
$\Delta{\boldsymbol{M}}$ $=-b_{e}^{s}\hat{v}_{f}^{2}({\boldsymbol{\mu}}_{e}{\boldsymbol{\mu}}_{e}^{\top}+{\boldsymbol{\nu}}_{e}{\boldsymbol{\nu}}_{e}^{\top})$ (18)
@@
$\left[{\boldsymbol{M}}_{c}+\Delta{\boldsymbol{M}}\right]^{-1}={\boldsymbol{M}}_{c}^{-1}+{\boldsymbol{M}}_{c}^{-1}\begin{pmatrix}{\boldsymbol{\mu}}_{e}&{\boldsymbol{\nu}}_{e}\end{pmatrix}$
$\quad\times\left[\frac{1}{b_{e}^{s}\hat{v}_{f}^{2}}{\boldsymbol{I}}-\begin{pmatrix}{\boldsymbol{\mu}}_{e}^{\top}\\ {\boldsymbol{\nu}}_{e}^{\top}\end{pmatrix}{\boldsymbol{M}}_{c}^{-1}\begin{pmatrix}{\boldsymbol{\mu}}_{e}&{\boldsymbol{\nu}}_{e}\end{pmatrix}\right]^{-1}\begin{pmatrix}{\boldsymbol{\mu}}_{e}^{\top}\\ {\boldsymbol{\nu}}_{e}^{\top}\end{pmatrix}{\boldsymbol{M}}_{c}^{-1}.$
@@
$({\boldsymbol{I}}-{\boldsymbol{M}}_{c}^{-1}{\boldsymbol{M}}_{o}){\boldsymbol{\mu}}_{e}=-{\boldsymbol{M}}_{c}^{-1}\,\Delta{\boldsymbol{M}}{\boldsymbol{\mu}}_{e}=-b_{e}^{s}\hat{v}_{f}^{2}{\boldsymbol{M}}_{c}^{-1}{\boldsymbol{\mu}}_{e}$
and an equivalent relation for ${\boldsymbol{\nu}}_{e}$. Using these relation as well as their transposes, we obtain
$\left[\frac{1}{b_{e}^{s}\,\hat{v}_{f}^{2}}{\boldsymbol{I}}-\begin{pmatrix}{\boldsymbol{\mu}}_{e}^{\top}\\ {\boldsymbol{\nu}}_{e}^{\top}\end{pmatrix}{\boldsymbol{M}}_{c}^{-1}\begin{pmatrix}{\boldsymbol{\mu}}_{e}&{\boldsymbol{\nu}}_{e}\end{pmatrix}\right$] (19)
@@
${\boldsymbol{M}}_{o}^{-1}$ $={\boldsymbol{M}}_{c}^{-1}+\left({\boldsymbol{I}}-{\boldsymbol{M}}_{c}^{-1}{\boldsymbol{M}}_{o}\right)\begin{pmatrix}{\boldsymbol{\mu}}_{e}&{\boldsymbol{\nu}}_{e}\end{pmatrix}$
@@
$\quad\times\begin{pmatrix}{\boldsymbol{\mu}}_{e}^{\top}\\ {\boldsymbol{\nu}}_{e}^{\top}\end{pmatrix}\left({\boldsymbol{I}}-{\boldsymbol{M}}_{o}{\boldsymbol{M}}_{c}^{-1}\right).$ (20)